A) If = -1.50 for a good, and price of the good decreases by 20 percent,...

A) If download?verifier=8ueeZh8AdM1ZknIbEinM8a = -1.50 for a good, and price of the good decreases by 20 percent, then:

Group of answer choices

quantity demanded will increase by 3 percent.

quantity demanded will decrease by 30 percent.

quantity demanded will increase by 30 percent.

total revenue will remain unchanged.

total revenue will decrease.

B)

The table below shows the total utility associated with various units of goods X and Y.
Table 7.5

Quantity

Total Utility from X

Total Utility from Y

1

24

85

2

42

130

3

56

160

4

66

185

5

74

200

6

80

210

7

84

215

Refer to Table 7.5. Given that the price of good X is $2 per unit, the price of good Y is $5 per unit, and that the consumer spends a total of $14 on both goods, what is the total utility associated with the utility-maximizing combination?

Group of answer choices

215

Indeterminate

154

172

84
Answer #1

(A) The price of good decreases by 20 percent.

It means % change in price = -20.

ed = (% change in quantity demanded / % change in price)

-1.50 = (% change in quantity demanded / -20)

% change in quantity demanded = (-1.5) (-20)

% change in quantity demanded = 30

Hence, the quantity demanded will increase by 30%.

Answer: Option (C).

(B)

Quantity TU from X MU from X MU from X / Price of X TU from Y MU from Y MU from Y / Price of Y
1 24 24 12 85 85 17
2 42 18 9 130 45 9
3 56 14 7 160 30 6
4 66 10 5 185 25 5
5 74 8 4 200 15 3
6 80 6 3 210 10 2
7 84 4 2 215 5 1

MU from X = Change in TU from X / Change in Quantity.

MU from Y = Change in TU from Y / Change in Quantity.

Utility maximized at the point where (MU from X / Price of X) = (MU from Y / Price of Y)

The above condition holds for four combinations of goods.

Combination 1: 2 units of X and 2 units of Y

Cost of buying this bundle = $2*2 + $5 * 2 =$14.

Budget to spend = $14.

hence this bundle is under the budget and holds the utility maximization condition. So, this would be the utility maximization bundle.

Combination 2: 4 units of X and 4 units of Y

Cost of buying this bundle = $2 * 4 + $5 * 4 = $28

Budget to spend = $14.

Hence, this bundle is not in the budget.

In a similar fashion, the other combinations of goods that satisfy the utility maximization condition are not in the budget except the combination 1.

Therefore, the utility-maximizing consumption bundle has 2 units if X and 2 units of Y.

Total utility from 2 units of X = 42

Total utility from 2 units of X = 130

Total utility associated with the utility-maximizing combination = 42 + 130

Total utility associated with the utility-maximizing combination = 172

Answer: Option (D) i.e., 172